OpenStudy (anonymous):

How do you find an equation of the line containing the given pair of points (-5,-4) and (-9,-1)

OpenStudy (anonymous):

The equation of a line is is the form of y=mx+b Where m is slope and b in the y intercept.

OpenStudy (anonymous):

First thing to do is find the slope! Slope=\[\frac{ y1-y2 }{ x1-x2 }\]

OpenStudy (anonymous):

After you find slope, you'd find the y-intercept by plugging in a pair of points into the equation. Example: (-5, -4) plugged into this equation : y = x + b -4 = (slope)(-5) + b